# The Contiguous Art Gallery Problem is in $\Theta(n\log n)$

### v1

@@ 1  Introduction
The Art Gallery problem is a classical problem in computational geometry. Given a simple polygon $P$ with $n$ vertices, the task is to compute the smallest set of guards such that every point $p\in P$ is visible to at least one guard. The problem was first posed by Victor Klee at a meeting of the Mathematical Association of America in 1973 and later formalized by Chvátal [17], who proved that $\lfloor n/3\rfloor$ guards always suffice and that this combinatorial bound is tight.
Since its introduction, the problem has been extensively studied. O’Rourke and Supowit [35] showed that the problem is NP-hard if $P$ is allowed to have holes, and Lee and Lin [30] strengthened this result by proving NP-hardness even when $P$ is a simple polygon without holes. Eidenbenz, Stamm, and Widmayer [21] established APX-hardness, and Bonnet and Miltzow [10] showed that the problem is W[1]-hard when parametrized by the number of guards. Finally, Abrahamsen, Adamaszek, and Miltzow [2] proved that the problem is $\exists\mathbb{R}$-complete.
@@
The first family of variants restricts the structure of the polygon $P$. Krohn and Nilsson [28] showed that the problem remains NP-hard when $P$ is $xy$-monotone. Schuchardt and Hecker [39] established NP-hardness for orthogonal polygons, and Tomás [42] extended this to the case of thin orthogonal polygons. Eidenbenz, Stamm, and Widmayer [21] proved NP-hardness even when $P$ is restricted to be a 1.5D terrain (an $x$-monotone polygonal chain that is closed by a base edge).
A second line of work constrains where guards may be placed. Lee and Lin [30] already proved NP-hardness when guards are restricted to the polygon vertices. Rieck and Scheffer [37] studied a dispersive variant where guards must maintain a minimum distance from one another, and showed NP-hardness as well. Related to this paradigm is the watchman problem, in which one or more guards move along paths inside the polygon. This problem also admits several sub-variants, the majority of which are NP-hard [16, 19, 34]. Those that can be solved in polynomial time require high-polynomial algorithms, with the best known running times of $O(n^{4})$ [15, 41, 12].
A third family of variants modifies the notion of visibility. Bhattacharya, Kumar, and Bodhayan [5] considered edge guards, where a point of $P$ is visible if it can be seen from some point on a guarding edge. Mahdavi, Seddighin, and Ghodsi [32] studied edge guards for rectilinear polygons in which new guarding edges may be introduced in $P$. Biedl and Mehrabi [7] analyzed rectilinear visibility, where two points see each other if there exists a rectilinear path between them contained in $P$. All of these variants were shown to be NP-hard. Worman and Keil [44] studied the same rectilinear visibility model but restricted $P$ to be rectilinear also; they obtained an $\tilde{O}(n^{17})$ polynomial-time algorithm. Hoorfar and Bagheri [26] considered rectangle vision, where two points $p,q\in P$ see each other if there exists a rectangle contained in $P$ having $p$ and $q$ as opposite corners. This problem can be solved in $\tilde{O}(n^{17})$ time for simple polygons but becomes NP-hard when holes are allowed [26]. Hillberg, Krohn, and Pahlow [25] showed that Art Gallery remains NP-hard when guards can only see to the right. Vaezi, Roy, and Ghodsi [43] studied visibility with mirrors that allow guards to see around corners, proving that several of these variants are also NP-hard.
@@
Only a fewArt Gallery variants are solvable in low-polynomial time, and all known results rely on very strong structural restrictions. Lee and Preparata [31] gave a linear-time algorithm to decide whether a polygon can be guarded by a single guard. Hoorfar and Bagheri [27] presented a linear-time algorithm for rectangle vision when $P$ is a histogram. Palios and Tzimas [36] restrict each guard’s visibility region to an orthogonal $r$-star and obtained a near-linear time algorithm. Biedl and Mehrabi [6] combined several restrictions—$P$ must be orthogonal [39], thin [42], and vision rectilinear [7]—and achieved a linear-time solution. Chen, Estivill-Castro, and Urrutia [14] gave a linear-time algorithm for 1.5D terrains when guards can only see to the right.
@@
Among the many known variants, some are motivated by geometric simplicity or mathematical curiosity, while others arise naturally from physical considerations. Laurentini [29] suggested guarding only the boundary of the polygon, based on the observation that in most art galleries, artworks are displayed on walls. This leads to the question: ‘Can we compute a smallest set of guards such that every point on $\partial P$ is visible to at least one guard?’
@@
Laurentini observed that this problem is NP-hard, and Stade [40] recently proved that it is $\exists\mathbb{R}$-complete. Thomas Shermer proposed the closely-related Contiguous Art Gallery problem. Here, each guard $g$ can only guard a contiguous segment $[u,v]$ of $\partial P$. The goal is to compute the smallest set of such contiguous guards whose visibility intervals together cover $\partial P$ (see Figure 1). At SoCG 2025, three independent works presented realRAM algorithms for this problem, each with comparable high-polynomial running times but based on entirely different methods. Let $k$ denote the size of the optimal solution. Biniaz et al. [8] analyzed the combinatorial structure of contiguous guards. They constructed a candidate set $\mathbb{C}$ of $O(n^{4})$ guards and guarantee that there exists an optimal solution that uses guards only from $\mathbb{C}$. They show how to compute $\mathbb{C}$, and combine these guards into an optimal solution, in $O(kn^{5}\log n)$ total time.
Merrild, Rysgaard, Schou, and Svenning [33] gave an $O(kn^{5}\log n)$-time greedy algorithm. For any point $u\in\partial P$, let $\textnormal{{next}}(u)$ denote the farthest point $v$ along $\partial P$ such that there exists a guard $(g,[u,v])$. Starting from $u$, one can recursively apply next until the resulting set $\{[u,v]\}$ covers $\partial P$. Such a sequence defines a revolution, which they show yields a solution of size at most $k+1$. They compute a revolution in $O(n^{2}\log n)$ time and prove that after $O(kn^{3})$ recursive revolutions, one must obtain an optimal solution, leading to an overall $O(kn^{5}\log n)$-time algorithm.
Robson, Spalding-Jamieson, and Zheng [38] took an analytical approach based on the same $\textnormal{{next}}(u)$ function. They parametrize the boundary $\partial P$ by $[1,2n+1)$, where points $i$ and $i+n$ correspond to the same vertex of $P$. Then we can treat next as a function from $[1,n+1)$ to $[1,2n+1)$. They proved that $\partial P$ can be partitioned into $O(n^{3})$ intervals, where next is of constant-complexity in each interval. They extend this to an $O(n^{6}\log n)$-time algorithm.
A subsequent joint version [9] presented an overview of these three approaches but did not provide a unified algorithm. Thus, there exist three independent approaches that use high polynomial time.
@@
Biniaz et al. [8] constructed a candidate set $\mathbb{C}$ of $O(n^{4})$ contiguous guards such that there exists an optimal solution that uses only guards from $\mathbb{C}$. We show that this condition is overly strong: it suffices that at least one guard from $\mathbb{C}$ occurs in an optimal solution. We compute in $O(n\log n)$ time a linear-size set $X$ of points $u\in\partial P$ such that there exists an optimal solution containing a contiguous guard $(g,[u,v])$. Running the revolution algorithm of Merrild et al. [33] for each $x\in X$ would then yield an optimal solution in $O(n^{3}\log n)$ time. However, their method effectively enumerates all interval pairs of a certain arrangement in $P$, leading to the high polynomial factor.
We avoid this enumeration by showing how to compute all values $\textnormal{{next}}(x)$ for $x\in X$ in $O(n\log n)$ total time using a second sweep around $\partial P$. Applying this second sweepline algorithm $k$ times yields an $O(kn\log n)$-time algorithm. To remove the remaining factor of $k$, we compute the revolutions for all $x\in X$ implicitly. To achieve this, we use the algebraic view of the next function introduced by Robson, Spalding-Jamieson, and Zheng [38], who showed that next can be represented as a piecewise function with $O(n^{3})$ constant-complexity pieces. Their approach constructs these pieces explicitly in $O(n^{3})$ time by partitioning $\partial P$ into intervals where next is defined by three vertices.
We revisit this construction and show that the number of distinct function pieces is in fact much smaller. This part of the paper is the most technical: we prove that the next function can be represented as a piecewise function with only $O(n)$ constant-complexity pieces, and that this representation can be computed in $O(n\log n)$ total time. Using a lazy segment tree evaluation technique due to Aggarwal et al. [4], we can then evaluate all revolutions from $x\in X$ simultaneously and obtain the optimal solution in $O(n\log n)$ time using $O(n)$ space. A particular point of emphasis for our approach is that the running time uses a single logarithmic factor. We show that our algorithms must maintain a convex hull under several types of updates. The best known general-purpose dynamic convex hull structure supports updates in $O(\log^{2}n)$ time. We can get by with an existing first-in-first-out structure to avoid any additional log-factors.
@@ 2  Preliminaries
We henceforth assume that $\partial P$ cannot be guarded by a single guard.
@@ 3  Technical Overview
Our primary contribution is a tight analysis of the Contiguous Art Gallery problem in the realRAM-model. We present an $O(n\log n)$-time algorithm, achieving a speedup by a factor of $\tilde{O}(kn^{4})$ compared to previous works [9, 33, 8, 38]. We complement this by a relatively simple lower bound of $\Omega(n\log n)$ in the realRAM-model.
Our algorithmic improvement arises from three key ideas. For any $x\in[1,n+1)$, let $\textnormal{{next}}(x)$ denote the maximal value $v$ such that there exists a guard $(g,[u,v])$. Our first idea is to construct a set $X$ of points along $\partial P$ with the property that there exists some $x\in X$ for which the greedy algorithm—obtained by recursively applying $\textnormal{{next}}(\cdot)$—yields an optimal solution. We construct a set $X$ of linear size and compute it in $O(n\log n)$ time. This result alone improves the running time of [9, 33] to $O(n^{3}\log n)$, as it suffices to compute what they call a revolution for each $x\in X$.
@@
Our final and most technical idea refines this approach. We partition $\partial P$ into intervals such that, for each interval $x_{1},x_{2}]$, the function $\textnormal{{next}}(x)$ on the domain $x_{1},x_{2}]$ can be described by a constant-complexity function. This observation was present in [9, 38], where they use $O(n^{3})$ intervals which they compute in near-cubic time. We show that $O(n)$ such intervals suffice, and compute them in $O(n\log n)$ time. We can combine these functions with our sliding-window procedure to create an algorithm that computes, for all $x\in X$, their full greedy sequences under the next function in $O(n\log n)$ total time—producing an optimal solution to the Contiguous Art Gallery problem.
@@ 3.1  Upper bound
We begin with a concept that appears in several previous approaches [9, 33, 8, 38]. For any $u\in\partial P$, let $v\in[u,2n+1)$ be the maximal value such that there exists a guard $(g,[u,v])$. This defines a function $\textnormal{{next}}:\partial P\to\partial P$ mapping $u$ to $v$. Let $k$ denote the minimum number of guards required to cover $\partial P$. If we choose any point $x\in\partial P$ and apply next a total of $k+1$ times, the resulting point lies beyond $x+n$. Our goal is to compute a set $X\subset\partial P$ such that for at least one $x\in X$ there exists an optimal solution of $k$ guards containing a guard $(g,[x,v])$. If we recursively apply the next function $k$ times to all points in $X$, this yields an optimal solution using $k$ guards. Hence, if $X$ has linear size and if we can evaluate next in $O(\log n)$ time, we obtain an $O(kn\log n)$-time algorithm.
@@
Let $\sigma$ be a conforming sliding sequence. Each bad dominator has a defining index pair $(i,j)$. Since reduced bad dominators are not strictly dominated by any other guard, every guard in $\mathcal{B}$ has its corresponding index pair $(i,j)\in\sigma$. From this observation we upper bound $|\mathcal{B}|$. Consider iterating over all $(i,j)\in\sigma$. By point-line duality, maintaining the convex visibility core $\mathcal{E}[i-1,j+1]$ during this iteration corresponds to maintaining the convex hull of a point set under first-in–first-out updates. Chan, Hershberger, and Pratt [13] show that the convex hull of such an update sequence of length $n$ has $O(n)$ vertices. This implies an algorithm to dynamically maintain $\mathcal{E}[i-1,j+1]$ subject to incrementing $i$ and $j$, and that $|\mathcal{B}|\in O(n)$. It remains to compute such a $\sigma$, and derive the sets $\mathcal{D}$ and $\mathcal{B}$, in $O(n\log n)$ time.
@@
For all subsequent algorithms, we use a conforming sliding sequence $\sigma$, which we compute in an online fashion starting from $(i,j)=(1,1)$ and repeatedly incrementing either $i$ or $j$. The choice of which index to increment follows a simple rule: while there exists a guard $(g,[u,v])$ such that $i,j+1]\subset[u,v]$, we increment $j$ and add the new pair $(i,j)$ to $\sigma$; otherwise, we increment $i$ and add the new pair $(i,j)$ to $\sigma$. By the result of [13], we can maintain $\mathcal{E}[i-1,j+2]$ as a balanced binary tree while incrementing $i$ and $j$. We show that, given $i$, $j$, and $\mathcal{E}[i-1,j+2]$, this decision can be made in $O(\log n)$ time, yielding an overall $O(n\log n)$ algorithm for constructing $\sigma$. From $\sigma$, we derive a linear-size set $\mathcal{B}_{\sigma}$ containing $\mathcal{B}$. We store the visibility of guards in $\mathcal{D}^{\prime}$ and $\mathcal{B}_{\sigma}$ in an segment tree $T$. Thus, we may obtain for any point $x\in\partial P$ the guard $(g,[u,v])\in\mathcal{D}^{\prime}\cup\mathcal{B}_{\sigma}$ that has $x\in[u,v]$ and maximizes $v$.
@@
In Section˜7, we consider the set $X$ and iterate over all $(i,j)\in\sigma$. For each $x\in X$, we perform a stabbing query on $T$ to find the maximum value $r$ such that there exists a guard $(g,[x,r])\in\mathcal{D}^{\prime}\cup\mathcal{B}_{\sigma}$. The value $\textnormal{{next}}(x)$ differs from $r$ only if it is realised by an ugly dominator. The value $\lfloor r\rfloor$ serves as a lower bound for the index $j$ such that $\textnormal{{next}}(x)\in(j,j+1]$, and by definition $(i,j)\in\sigma$. These observations yield an algorithm that iterates over $\sigma$ and computes $\textnormal{{next}}(x)$ for all $x\in X$ in $O(n\log n)$ total time. Repeating this process $k$ times produces an optimal solution in $O(kn\log n)$ time. This is already a major improvement over the state-of-the-art [9, 33, 8, 38] and serves as a crucial subroutine for our $O(n\log n)$-time algorithm.
@@
We compute a representation of next as a piecewise, possibly degenerate, Möbius transformation. This approach was first proposed in [9, 38], where the resulting function consisted of $O(n^{3})$ pieces and was computed in $\tilde{O}(n^{5})$ time. In contrast, we use our dominators to define a function with only $O(n)$ pieces, which can be computed in $O(n\log n)$ time. Specifically, in Section˜8, we partition the boundary $\partial P$ into intervals $I$ such that either:
@@
We first prove that the vertices of the shortest path from $x$ to $\textnormal{{next}}(x)$ behave in a structured way, unless the optimal solution consists of at most $3$ guards. These vertices play a key role in defining the ugly dominator $g_{u}$, and the resulting structural properties guarantee both the existence and computability of the $O(n)$ Möbius transformations in $O(n\log n)$ total time. Finally, in Section˜9, we combine all components to obtain our $O(n\log n)$ algorithm. We apply our $O(kn\log n)$ algorithm to check whether a solution with $k\leq 3$ guards exists, in $O(n\log n)$ time. If not, we use the techniques of Aggarwal et al. [4] to compute the smallest $k$ such that there exists a point $x\in X$ for which $\textnormal{{next}}^{k}(x)\geq x+n$, or conversely, $x,\textnormal{{next}}^{k}(x)]=\partial P$, where $\textnormal{{next}}^{k}(\cdot)$ is the composition of next with itself $k$ times, all within $O(n\log n)$ time.

### v2

@@ 1  Introduction
The Art Gallery problem is a classical problem in computational geometry. Given a simple polygon $P$ with $n$ vertices, the task is to compute the smallest set of guards such that every point $p\in P$ is visible to at least one guard. The problem was first posed by Victor Klee at a meeting of the Mathematical Association of America in 1973 and later formalized by Chvátal [chvatal1975combinatorial], who proved that $\lfloor n/3\rfloor$ guards always suffice and that this combinatorial bound is tight.
Since its introduction, the problem has been extensively studied. O’Rourke and Supowit [Rourke1983hard] showed that the problem is NP-hard if $P$ is allowed to have holes, and Lee and Lin [lee1986VertexNPhard] strengthened this result by proving NP-hardness even when $P$ is a simple polygon without holes. Eidenbenz, Stamm, and Widmayer [EidenbenzStammWidmayer2001] established APX-hardness, and Bonnet and Miltzow [Bonnet2020Parametrized] showed that the problem is W[1]-hard when parametrized by the number of guards. Finally, Abrahamsen, Adamaszek, and Miltzow [abrahamsen2018ETRhard] proved that the problem is $\exists\mathbb{R}$-complete.
@@
The first family of variants restricts the structure of the polygon $P$. Krohn and Nilsson [KrohnNilsson2012] showed that the problem remains NP-hard when $P$ is $xy$-monotone. Schuchardt and Hecker [SchuchardtHecker1995] established NP-hardness for orthogonal polygons, and Tomás [Tomas2013ThinOrth] extended this to the case of thin orthogonal polygons. Eidenbenz, Stamm, and Widmayer [EidenbenzStammWidmayer2001] proved NP-hardness even when $P$ is restricted to be a 1.5D terrain (an $x$-monotone polygonal chain that is closed by a base edge).
A second line of work constrains where guards may be placed. Lee and Lin [lee1986VertexNPhard] already proved NP-hardness when guards are restricted to the polygon vertices. Rieck and Scheffer [Rieck2024Dispersive] studied a dispersive variant where guards must maintain a minimum distance from one another, and showed NP-hardness as well. Related to this paradigm is the watchman problem, in which one or more guards move along paths inside the polygon. This problem also admits several sub-variants, the majority of which are NP-hard [ChinNtafos1988, DumitrescuToth2012, Mitchell2013]. Those that can be solved in polynomial time require high-polynomial algorithms, with the best known running times of $O(n^{4})$ [ChinNtafos1991, Tan2001, CarlssonJonssonNilsson1999].
A third family of variants modifies the notion of visibility. Bhattacharya, Kumar, and Bodhayan [BHATTACHARYA2017109] considered edge guards, where a point of $P$ is visible if it can be seen from some point on a guarding edge. Mahdavi, Seddighin, and Ghodsi [MAHDAVI2020163] studied edge guards for rectilinear polygons in which new guarding edges may be introduced in $P$. Biedl and Mehrabi [Biedl2017Mehrabi] analyzed rectilinear visibility, where two points see each other if there exists a rectilinear path between them contained in $P$. All of these variants were shown to be NP-hard. Worman and Keil [Worman2007Decomposition] studied the same rectilinear visibility model but restricted $P$ to be rectilinear also; they obtained an $\tilde{O}(n^{17})$ polynomial-time algorithm. Hoorfar and Bagheri [hoorfar2017guarding] considered rectangle vision, where two points $p,q\in P$ see each other if there exists a rectangle contained in $P$ having $p$ and $q$ as opposite corners. This problem can be solved in $\tilde{O}(n^{17})$ time for simple polygons but becomes NP-hard when holes are allowed [hoorfar2017guarding]. Hillberg, Krohn, and Pahlow [Hillberg2022complexity] showed that Art Gallery remains NP-hard when guards can only see to the right. Vaezi, Roy, and Ghodsi [Vaezi2023Reflective] studied visibility with mirrors that allow guards to see around corners, proving that several of these variants are also NP-hard.
@@
Only a few Art Gallery variants are solvable in low-polynomial time, and all known results rely on very strong structural restrictions. Lee and Preparata [Lee1979ptimal] gave a linear-time algorithm to decide whether a polygon can be guarded by a single guard. Hoorfar and Bagheri [HoorfarBagheri2017HiddenGuards] presented a linear-time algorithm for rectangle vision when $P$ is a histogram. Palios and Tzimas [Palios2013-star-cover] restrict each guard’s visibility region to an orthogonal $r$-star and obtained a near-linear time algorithm. Biedl and Mehrabi [BiedlMehrabi2016_rGuardingThinOrthogonal] combined several restrictions—$P$ must be orthogonal [SchuchardtHecker1995], thin [Tomas2013ThinOrth], and vision rectilinear [Biedl2017Mehrabi]—and achieved a linear-time solution. Chen, Estivill-Castro, and Urrutia [ChenEstivillCastroUrrutia1995] gave a linear-time algorithm for 1.5D terrains when guards can only see to the right.
@@
Among the many known variants, some are motivated by geometric simplicity or mathematical curiosity, while others arise naturally from physical considerations. Laurentini [Laurentini1999GuardingWalls] suggested guarding only the boundary of the polygon, based on the observation that in most art galleries, artworks are displayed on walls. This leads to the question: ‘Can we compute a smallest set of guards such that every point on $\partial P$ is visible to at least one guard?’
@@
Laurentini observed that this problem is NP-hard, and Stade [stade2025interiorBoundaryHard] recently proved that it is $\exists\mathbb{R}$-complete. Thomas Shermer proposed the closely-related Contiguous Art Gallery problem. Here, each guard $g$ can only guard a contiguous segment $[u,v]$ of $\partial P$. The goal is to compute the smallest set of such contiguous guards whose visibility intervals together cover $\partial P$ (see Figure 1). At SoCG 2025, three independent submissions presented realRAM algorithms for this problem, each with comparable high-polynomial running times but based on entirely different methods. Let $k$ denote the size of the optimal solution. Biniaz et al. [biniaz2024contiguous] analyzed the combinatorial structure of contiguous guards. They constructed a candidate set $\mathbb{C}$ of $O(n^{4})$ guards and guarantee that there exists an optimal solution that uses one of the guards from $\mathbb{C}$. They show how to compute $\mathbb{C}$, and how to compute from these an optimal solution n $O(kn^{5}\log n)$ total time.
Merrild, Rysgaard, Schou, and Svenning [merrild2024contiguous] gave an $O(kn^{5}\log n)$-time greedy algorithm. For any point $u\in\partial P$, let $\textnormal{{next}}(u)$ denote the farthest point $v$ along $\partial P$ such that there exists a guard $(g,[u,v])$. Starting from $u$, one can recursively apply next until the resulting set $\{[u,v]\}$ covers $\partial P$. Such a sequence defines a revolution, which they show yields a solution of size at most $k+1$. They compute a revolution in $O(n^{2}\log n)$ time and prove that after $O(kn^{3})$ recursive revolutions, one must obtain an optimal solution, leading to an overall $O(kn^{5}\log n)$-time algorithm.
Robson, Spalding-Jamieson, and Zheng [RobsonSpaldingZheng2024_AnalyticArcCover] took an analytical approach based on the same $\textnormal{{next}}(u)$ function. They parametrize the boundary $\partial P$ by $[1,2n+1)$, where points $i$ and $i+n$ correspond to the same vertex of $P$. Then we can treat next as a function from $[1,n+1)$ to $[1,2n+1)$. They proved that $\partial P$ can be partitioned into $O(n^{3})$ intervals, where next is of constant-complexity in each interval. They extend this to an $O(n^{6}\log n)$-time algorithm.
A subsequent joint version [theEnemy] presented an overview of these three approaches but did not provide a unified algorithm. Thus, there exist three independent approaches that use high polynomial time.
@@
Biniaz et al. [biniaz2024contiguous] constructed a candidate set $\mathbb{C}$ of $O(n^{4})$ contiguous guards such that there exists an optimal solution that uses one guard from $\mathbb{C}$. We show that a linear-size set suffices. We compute in $O(n\log n)$ time a linear-size set $X$ of points $u\in\partial P$ such that there exists an optimal solution containing a contiguous guard $(g,[u,v])$. Running the revolution algorithm of Merrild et al. [merrild2024contiguous] for each $x\in X$ would then yield an optimal solution in $O(n^{3}\log n)$ time. However, their method effectively enumerates all interval pairs of a certain arrangement in $P$, leading to the high polynomial factor.
We avoid this enumeration by showing how to compute all values $\textnormal{{next}}(x)$ for $x\in X$ in $O(n\log n)$ total time using a second sweep around $\partial P$. Applying this second sweepline algorithm $k$ times yields an $O(kn\log n)$-time algorithm. To remove the remaining factor of $k$, we compute the revolutions for all $x\in X$ implicitly. To achieve this, we use the algebraic view of the next function introduced by Robson, Spalding-Jamieson, and Zheng [RobsonSpaldingZheng2024_AnalyticArcCover], who showed that next can be represented as a piecewise function with $O(n^{3})$ constant-complexity pieces. Their approach constructs these pieces explicitly in $O(n^{3})$ time by partitioning $\partial P$ into intervals where next is defined by three vertices.
We revisit this construction and show that the number of distinct function pieces is in fact much smaller. This part of the paper is the most technical: we prove that the next function can be represented as a piecewise function with only $O(n)$ constant-complexity pieces, and that this representation can be computed in $O(n\log n)$ total time. Using a lazy segment tree evaluation technique due to Aggarwal et al. [aggarwal1989finding], we can then evaluate all revolutions from $x\in X$ simultaneously and obtain the optimal solution in $O(n\log n)$ time using $O(n)$ space. A particular point of emphasis for our approach is that the running time uses a single logarithmic factor. We show that our algorithms must maintain a convex hull under several types of updates. The best known general-purpose dynamic convex hull structure supports updates in $O(\log^{2}n)$ time. We can get by with an existing first-in-first-out structure to avoid any additional log-factors.
@@ 3  Technical Overview
Our primary contribution is a tight analysis of the Contiguous Art Gallery problem in the realRAM-model. We present an $O(n\log n)$-time algorithm, achieving a speedup by a factor of $\tilde{O}(kn^{4})$ compared to previous works [theEnemy, merrild2024contiguous, biniaz2024contiguous, RobsonSpaldingZheng2024_AnalyticArcCover]. We complement this by a relatively simple lower bound of $\Omega(n\log n)$ in the realRAM-model.
Our algorithmic improvement arises from three key ideas. For any $x\in[1,n+1)$, let $\textnormal{{next}}(x)$ denote the maximal value $v$ such that there exists a guard $(g,[u,v])$. Our first idea is to construct a set $X$ of points along $\partial P$ with the property that there exists some $x\in X$ for which the greedy algorithm—obtained by recursively applying $\textnormal{{next}}(\cdot)$—yields an optimal solution. We construct a set $X$ of linear size and compute it in $O(n\log n)$ time. This result alone improves the running time of [theEnemy, merrild2024contiguous] to $O(n^{3}\log n)$, as it suffices to compute what they call a revolution for each $x\in X$.
@@
Our final and most technical idea refines this approach. We partition $\partial P$ into intervals such that, for each interval $x_{1},x_{2}]$, the function $\textnormal{{next}}(x)$ on the domain $x_{1},x_{2}]$ can be described by a constant-complexity function. This observation was present in [theEnemy, RobsonSpaldingZheng2024_AnalyticArcCover], where they use $O(n^{3})$ intervals which they compute in near-cubic time. We show that $O(n)$ such intervals suffice, and compute them in $O(n\log n)$ time. We can combine these functions with our sliding-window procedure to create an algorithm that computes, for all $x\in X$, their full greedy sequences under the next function in $O(n\log n)$ total time—producing an optimal solution to the Contiguous Art Gallery problem.
@@ 3.1  Upper bound
We begin with a concept that appears in several previous approaches [theEnemy, merrild2024contiguous, biniaz2024contiguous, RobsonSpaldingZheng2024_AnalyticArcCover]. For any $u\in\partial P$, let $v\in[u,2n+1)$ be the maximal value such that there exists a guard $(g,[u,v])$. This defines a function $\textnormal{{next}}:\partial P\to\partial P$ mapping $u$ to $v$. Let $k$ denote the minimum number of guards required to cover $\partial P$. If we choose any point $x\in\partial P$ and apply next a total of $k+1$ times, the resulting point lies beyond $x+n$. Our goal is to compute a set $X\subset\partial P$ such that for at least one $x\in X$ there exists an optimal solution of $k$ guards containing a guard $(g,[x,v])$. If we recursively apply the next function $k$ times to all points in $X$, this yields an optimal solution using $k$ guards. Hence, if $X$ has linear size and if we can evaluate next in $O(\log n)$ time, we obtain an $O(kn\log n)$-time algorithm.
@@
Let $\sigma$ be a conforming sliding sequence. Each bad dominator has a defining index pair $(i,j)$. Since reduced bad dominators are not strictly dominated by any other guard, every guard in $\mathcal{B}$ has its corresponding index pair $(i,j)\in\sigma$. From this observation we upper bound $|\mathcal{B}|$. Consider iterating over all $(i,j)\in\sigma$. By point-line duality, maintaining the convex visibility core $\mathcal{E}[i-1,j+1]$ during this iteration corresponds to maintaining the convex hull of a point set under first-in–first-out updates. Chan, Hershberger, and Pratt [Chan2019Time-Windowed] show that the convex hull of such an update sequence of length $n$ has $O(n)$ vertices. This implies an algorithm to dynamically maintain $\mathcal{E}[i-1,j+1]$ subject to incrementing $i$ and $j$, and that $|\mathcal{B}|\in O(n)$. It remains to compute such a $\sigma$, and derive the sets $\mathcal{D}$ and $\mathcal{B}$, in $O(n\log n)$ time.
@@
For all subsequent algorithms, we use a conforming sliding sequence $\sigma$, which we compute in an online fashion starting from $(i,j)=(1,1)$ and repeatedly incrementing either $i$ or $j$. The choice of which index to increment follows a simple rule: while there exists a guard $(g,[u,v])$ such that $i,j+1]\subset[u,v]$, we increment $j$ and add the new pair $(i,j)$ to $\sigma$; otherwise, we increment $i$ and add the new pair $(i,j)$ to $\sigma$. By the result of [Chan2019Time-Windowed], we can maintain $\mathcal{E}[i-1,j+2]$ as a balanced binary tree while incrementing $i$ and $j$. We show that, given $i$, $j$, and $\mathcal{E}[i-1,j+2]$, this decision can be made in $O(\log n)$ time, yielding an overall $O(n\log n)$ algorithm for constructing $\sigma$. From $\sigma$, we derive a linear-size set $\mathcal{B}_{\sigma}$ containing $\mathcal{B}$. We store the visibility of guards in $\mathcal{D}^{\prime}$ and $\mathcal{B}_{\sigma}$ in an segment tree $T$. Thus, we may obtain for any point $x\in\partial P$ the guard $(g,[u,v])\in\mathcal{D}^{\prime}\cup\mathcal{B}_{\sigma}$ that has $x\in[u,v]$ and maximizes $v$.
@@
In Section˜7, we consider the set $X$ and iterate over all $(i,j)\in\sigma$. For each $x\in X$, we perform a stabbing query on $T$ to find the maximum value $r$ such that there exists a guard $(g,[x,r])\in\mathcal{D}^{\prime}\cup\mathcal{B}_{\sigma}$. The value $\textnormal{{next}}(x)$ differs from $r$ only if it is realised by an ugly dominator. The value $\lfloor r\rfloor$ serves as a lower bound for the index $j$ such that $\textnormal{{next}}(x)\in(j,j+1]$, and by definition $(i,j)\in\sigma$. These observations yield an algorithm that iterates over $\sigma$ and computes $\textnormal{{next}}(x)$ for all $x\in X$ in $O(n\log n)$ total time. Repeating this process $k$ times produces an optimal solution in $O(kn\log n)$ time. This is already a major improvement over the state-of-the-art [theEnemy, merrild2024contiguous, biniaz2024contiguous, RobsonSpaldingZheng2024_AnalyticArcCover] and serves as a crucial subroutine for our $O(n\log n)$-time algorithm.
@@
We compute a representation of next as a piecewise, possibly degenerate, Möbius transformation. This approach was first proposed in [theEnemy, RobsonSpaldingZheng2024_AnalyticArcCover], where the resulting function consisted of $O(n^{3})$ pieces and was computed in $\tilde{O}(n^{5})$ time. In contrast, we use our dominators to define a function with only $O(n)$ pieces, which can be computed in $O(n\log n)$ time. Specifically, in Section˜8, we partition the boundary $\partial P$ into intervals $I$ such that either:
@@
We first prove that the vertices of the shortest path from $x$ to $\textnormal{{next}}(x)$ behave in a structured way, unless the optimal solution consists of at most $3$ guards. These vertices play a key role in defining the ugly dominator $g_{u}$, and the resulting structural properties guarantee both the existence and computability of the $O(n)$ Möbius transformations in $O(n\log n)$ total time. Finally, in Section˜9, we combine all components to obtain our $O(n\log n)$ algorithm. We apply our $O(kn\log n)$ algorithm to check whether a solution with $k\leq 3$ guards exists, in $O(n\log n)$ time. If not, we use the techniques of Aggarwal et al. [aggarwal1989finding] to compute the smallest $k$ such that there exists a point $x\in X$ for which $\textnormal{{next}}^{k}(x)\geq x+n$, or conversely, $x,\textnormal{{next}}^{k}(x)]=\partial P$, where $\textnormal{{next}}^{k}(\cdot)$ is the composition of next with itself $k$ times, all within $O(n\log n)$ time.